OpenStudy (anonymous):

10(3x+1)=2(10x-5) wats x

OpenStudy (anonymous):

30x + 10 = 20x - 10 -20x -20x 10x + 10 = -10 10x = 0 x = 0

OpenStudy (kira_yamato):

Hope this helps

OpenStudy (kamille):

Nope, you arent corret kpi, because the answer should be -2

OpenStudy (kamille):

10x + 10 = -10 10x = 0 it it the mistake 10x=-10-10 10x=-20|:10 x=-2

OpenStudy (anonymous):

answer is -2.... 10(3x+1)=2(10x-5) 30x+10=20x-10 by solving it we get x=-2....!!!!!!!